# PTDiffusion: Free Lunch for Generating Optical Illusion Hidden Pictures with Phase-Transferred Diffusion Model

### v2

@@ 4.2 Quantitative results
For quantitative evaluations, we measure image visual quality using Aesthetic Score ($\uparrow$) predicted by the pre-trained LAION Aesthetics Predictor V2 model, and measure text fidelity using CLIP Score ($\uparrow$), i.e., the image-text cosine similarity. Considering that weak I2I appearance binding is preferred for generating optical illusion pictures, we measure model’s appearance modification ability by evaluating LPIPS ($\uparrow$) between the reference image and the generated illusion image. We evaluate on 80 reference images with each one paired with at least 3 text prompts. Results reported in Tab. 1 demonstrate that our method achieves leading performance for all the aforementioned metrics.
We conduct quantitative ablation studies about the influence of the decayed phase transfer, the refining stage, and the DDIM inversion w.r.t the generation visual quality and text fidelity. Results displayed in Tab. 2 show that the absence of any one of these ingredients results in declined aesthetic score and image-text CLIP similarity score, which is basically in line with the qualitative results displayed in Fig. 9 and Fig. 10. Moreover, missing the refining stage causes the most performance drop in these two metrics.
To quantitatively prove the hidden image perceptibility control ability of our proposed asynchronous phase transfer, we quantify hidden content visual prominence as structure similarity between the input and output image pair, for which we use DINO-ViT self-similarity score [44] as the metric. As reported in Tab. 3, the I2I structure similarity continuously grows with the increase of the async distance $d$, which tallies with the qualitative results of Fig. 11. It is also worth noting that the increase of $d$ does not lead to drastic drop in CLIP Score, indicating that our proposed asynchronous phase transfer promotes hidden content prominence without noticeably sacrificing text fidelity.
@@
For aesthetic assessment of the generated optical illusion hidden pictures, we resort to user study for subjective evaluation. Based on the unique visual characteristics of the illusion pictures, we invite 16 participants to score the generation results of different methods on a scale of 1-10 from the following two perspectives: (i) contextual naturalness, i.e., to what extent the reference image is naturally and reasonably blended into the target semantic content; (ii) illusion balance, i.e., well-balanced visual discernibility of the target semantic content and the original hidden content, rather than visual prominence of only one side. The average user ratings to all the compared methods in Tab. 1 are reported in Fig. 12, our method outscores other approaches by a large margin in both two dimensions, demonstrating significant advantage of our PTDiffusion in illusion picture synthesis.

### v3

@@ 4.2 Quantitative results
For quantitative evaluations, we measure image visual quality using Aesthetic Score ($\uparrow$) predicted by the pre-trained LAION Aesthetics Predictor V2 model, and measure text fidelity using CLIP Score ($\uparrow$), i.e., the image-text cosine similarity. Considering that weak I2I appearance binding is preferred for generating optical illusion pictures, we measure model’s appearance modification ability by evaluating LPIPS ($\uparrow$) between the reference image and the generated illusion image. We test on 80 reference images with each one paired with at least 3 text prompts. Results reported in Tab. 1 demonstrate that our method achieves leading performance for all the aforementioned metrics.
We conduct quantitative ablation studies about the influence of the phase transfer decay, the refining stage, and the DDIM inversion w.r.t the image generation quality and text fidelity. Results displayed in Tab. 2 show that missing any one of these ingredients results in declined aesthetic score and image-text CLIP similarity score, which is basically in line with the qualitative results displayed in Fig. 9 and Fig. 10. Moreover, the absence of the refining stage causes the most performance drop for these two metrics.
To quantitatively prove the hidden content perceptibility control ability of our proposed asynchronous phase transfer, we quantify hidden content visual prominence as structure similarity between the input and output image pair, for which we use DINO-ViT self-similarity score [44] as the metric. As reported in Tab. 3, the I2I structure similarity continuously grows with the increase of the async distance $d$, which tallies with the qualitative results of Fig. 11. It is also worth noting that the increase of $d$ does not lead to drastic drop in CLIP Score, indicating that our proposed asynchronous phase transfer promotes hidden content discernibility without noticeably sacrificing text fidelity.
@@
For aesthetic assessment of the generated optical illusion hidden pictures, we resort to user study for subjective evaluation. Based on the unique visual characteristics of illusion pictures, we invite 16 participants to score the generation results of different methods on a scale of 1-10 from the following two perspectives: (i) contextual naturalness, i.e., to what extent the reference image is naturally and reasonably blended into the textual scene content; (ii) illusion balance, i.e., well-balanced visual discernibility of the textual scene content and the original hidden content, rather than visual prominence of only one side. The average user ratings of all the compared methods in Tab. 1 are reported in Fig. 12, our method outscores other approaches by a large margin in both two dimensions, subjectively indicating significant advantage of our PTDiffusion in illusion picture synthesis.